OpenStudy (anonymous):

find dy/dx for y=1+3x-x2

OpenStudy (anonymous):

3-2x

OpenStudy (anonymous):

can you please explain how you have done this

OpenStudy (anonymous):

The power rule states that the derivative of an exponential, such as 2x^2 is the coefficient times the exponent, whilst the exponent is subtracted by one. So, in my exampled you'd get 4x. Constants equal to zeros. As in, 1+2x^2 would just be 4x. Apply that to your problem.

OpenStudy (anonymous):

i got your point... see i have also did same but my answer is coming 3-2=1 not 3-2x

OpenStudy (anonymous):

can you please elaborate it

OpenStudy (anonymous):

That last value in your poblem, is it x2 or x^2?

OpenStudy (anonymous):

x square

OpenStudy (anonymous):

So, the 1 becomes a zero. the 3x loses the x, and just imagine the x^2 looked like this: -1x^2, multiply 2 by the -1, and then subtract the 2 by 1. You get -2x.

OpenStudy (anonymous):

i am sending you the way i have done the problem please check it

OpenStudy (anonymous):



OpenStudy (anonymous):

Dude, forget the dx and dys. You're complicating yourself. Just multiply the exponents by their coefficients and subtract said exponents by one. That's it. x2 should be written x^2.

OpenStudy (anonymous):

5x^5-3x^3+2x+100 = 25x^4-9x^2+2 <--- see it?

OpenStudy (zzr0ck3r):

\(1+3x-x^2=18x^0+3x^1-x^2\) so \(\frac{d}{dx}(1+3x-x^2)=0*1x^{0-1}+1*3x^{1-1}-2x^{2-1}=0+3x^0-2x^1=3-2x\)

OpenStudy (zzr0ck3r):

that shuold say 1 not 18...

OpenStudy (anonymous):

@zz0c3k you have not used dy and dx???/

OpenStudy (anonymous):

dx and dy just refer to the derivative in respect to those. Unless you get into implicit differentiation, they are kind of useless when doing something like what you are doing.

OpenStudy (anonymous):

the question you gave me above i tried to solve it and got the ans 25x\[25x ^{4} - 9x ^{2} +2x+100\]

OpenStudy (anonymous):

Numbers without letters = 0

OpenStudy (anonymous):

Numbers with only one letters are just those numbers. 2x = 2.

OpenStudy (zzr0ck3r):

\(\frac{dy}{dx}\) is the same thing as \(f(x)\) is a sense if y = f(x) we have \(\frac{d(f(x))}{dx}\) so when we take the derivative we note this by \(\frac{d}{dx}(the \ function)=\frac{d}{dx}(f(x))\) example let \(y = f(x) = x^2\) then we write \(\frac{dy}{dx}=\frac{d}{dx}(x^2)=2x\)

OpenStudy (anonymous):

but my professor told that we cannot ignore x we have to take 2x

OpenStudy (anonymous):

So see it like this 2x^1. That would be 2(1)x^1-1 = 2x^0 and x^0 =1 so 2(1) = 2

OpenStudy (anonymous):

see i got your point can you please solve another question

OpenStudy (anonymous):

Why not try yourself?

OpenStudy (zzr0ck3r):

a much better idea\(\huge \uparrow\)

OpenStudy (anonymous):

i have tried bt i am getting same problem in the question the question says.. in each of the following case evaluate dy/dx when x=2 \[y=ax ^{2} +bx+c\]

OpenStudy (anonymous):

@zzr0ck3r Jfc good luck.

OpenStudy (anonymous):

please i need your help

OpenStudy (zzr0ck3r):

if x = 2, then you are asking to take the derivative of a constant, what is the derivative of a constant?

OpenStudy (anonymous):

i have typed what the question says we have to find dy/dx for the following equation and after finding we have to put the value of x

OpenStudy (zzr0ck3r):

ahh, well \(y'=2ax+b\) so \(4a+b\)

OpenStudy (anonymous):

yaaa its correct how can you do it so fast!!! please elaborate it

OpenStudy (anonymous):

practice and it becomes second nature

OpenStudy (zzr0ck3r):

ax^2+bx+c we move the 2 down and decrement it by 1 and get 2ax^1 = 2ax then we look at the bx term, the derivative is just the constant, so we have b so 2ax+b

OpenStudy (zzr0ck3r):

the c term goes to 0

OpenStudy (anonymous):

the derivative b is constant how?

OpenStudy (anonymous):

2ax.. but hows it possible as i think that dx and x gets cut and we are left with 2a only..please correct my point

OpenStudy (zzr0ck3r):

the derivative of bx is b for all b....

OpenStudy (zzr0ck3r):

I dont know what you mean by x and dx get cut, I think you are getting confused with notation don't even think of the notation as dy/dx right now, just think of the derivative of f(x) is f'(x)

OpenStudy (zzr0ck3r):

or explain your question more

OpenStudy (anonymous):

see my solution